# Out-of-equlibrium inference of feeding rates through population data from generic consumer-resource stochastic dynamics

### v2

@@ 3.2 Out-of-equilibrium probability distribution
This result agrees with those obtained at equilibrium, in the limit $t\to\infty$. As explained in the previous subsection, we also found a multinomial distribution in the stationary regime. However, the out-of-equilibrium distribution is not multinomial in general: note that the generating function for non-zero initial populations, Eq. (66), does not match the one one associated to a multinomial distribution, which is given by Eq. (67), unless there are no handling predators initially. $P_{\bm{n}}(t)$ follows a multinomial distribution only if $\bm{\hat{n}}=\bm{0}$. Irrespective of the set of initial conditions, however, the generating function (66) has the correct limit as $t\to\infty$. Indeed, if we take this limit on both sides of Eq. (66) we get
$\lim_{t\to\infty}G(\bm{z},t)=\left(\frac{\nu+R({\bm{z}})}{\nu+Q}\right)^{n_{A}^{0}}=\biggl{(}p_{A}+\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}},$ (72)
@@ 4 Effect of steady-state and out-of-equilibrium sampling
We replicate the experiment a number of times $T$, each replica simulating the stochastic dynamics (using Gillespie (1977) method) with $n_{A}^{0}$ free predators initially, and record abundances at time $t_{j}$ for each realization $j=1,\dots,T$. As shown in B, it is important that at least two sampling times are different (i.e., $t_{j}\neq t_{k}$ for some $j\neq k$). Otherwise the model is not identifiable. In practice, from an empirical point of view, it is very unlikely that independent system samples are obtained at precisely the same exact time.
@@
$M(\{t_{j}\})=\begin{pmatrix}n_{1,1}(t_{1})&n_{1,2}(t_{2})&\cdots&n_{1,T}(t_{T})\\ \vdots&\vdots&\ddots&\vdots\\ n_{S,1}(t_{1})&n_{S,2}(t_{2})&\cdots&n_{S,T}(t_{T})\end{pmatrix},$ (73)
where columns comprise vectors $\bm{n}_{j}=(n_{1,j},\dots,n_{S,j})$ reporting observed abundances of handling predators at sampling time $t_{j}$. Free predator abundance at time $t_{j}$ is computed with $n_{A}(t_{j})=n_{A}^{0}-\sum_{i=1}^{S}n_{i,j}(t_{j})$. If each column of $M(\{t_{j}\})$ is a sample of the Holling type II multi-resource dynamics, then the probability $P_{\bm{n}_{j}}(t_{j}|\bm{\theta})$, computed using Eqs. (33) and (68), is the probability of observing a population vector $(n_{A}(t_{j}),\bm{n}_{j})$ given a model parameter vector, $\bm{\theta}:=(\alpha,\nu)$ —notice that resource densities, $x_{R_{i}}=n_{R_{i}}/N$, the system volume, $N$, and the total number of consumers, $n_{A}^{0}$, are kept constant by assumption and therefore are not regarded as parameters in $\bm{\theta}$. Matrix $M(\{t_{j}\})$ gathers $T$ independent observations, so the function that assigns a linkelihood to parameter vector $\bm{\theta}$ given the data matrix $M(\{t_{j}\})$ is multiplicative:
$\mathcal{L}(\bm{\theta}|M(\{t_{j}\}))=\prod_{j=1}^{T}P_{\bm{n}_{j}}(t_{j}|\bm{\theta}).$ (74)
The log-likelihood function,
@@
Figure 1: Parameter estimation using equilibrium samples. We generated a single realization of the data matrix $M(\{t_{j}\})$, including $T=100$ independent model realizations starting from $n_{A}^{0}=100$ searching predators at $t=0$, using Gillespie’s stochastic simulation algorithm up to $T=100$ different ending times uniformly drawn in the interval $[50\tau,100\tau]$ to ensure stationarity, with $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ the characteristic time of the dynamics. True rates are $\alpha=5$, $\nu=10$, and individual consumers feed on $S=5$ resources with densities $\bm{x_{R}}=(0.9,0.7,0.5,0.3,0.1)$. Left: for that synthetic dataset, we run a L-BFGS-B optimization algorithm for the log-likelihood function (75) starting from different initial conditions, and we recorded the values of $(\alpha^{\star},\nu^{\star})$ at which the maximum was reached. Optimal parameter values follow a perfect straight line. The distributions of $\alpha^{\star}$ and $\nu^{\star}$ are compatible with uniform distributions, and the optimal values of $\log\mathcal{L}$ keep constant up to round-off errors. Overall, this indicates a line of infinite optima. Right: we replicated the previous experiment 200 times and, for each replica, we estimated the ratio $\nu/\alpha$ by fitting the slope of the straight line. We obtained the distribution in the right panel, which is compatible with a Gaussian with mean 2 and variance extracted from data. The vertical, dashed line indicates the true ratio.
@@
Now, consider the effect of sampling times and assume that matrix $M(\{t_{j}\})$ is obtained when $t_{j}$ are “large enough” compared to the characteristic time of the dynamics, which is given by $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ —see Eq. (68). In Fig. 1, in silico data (i.e., abundances of each behavioral type) were generated by stochastic simulation of a system with $S$ resources (with fixed densities $x_{R_{i}}$) consumed by $n_{A}^{0}$ predator individuals. A number $T$ of independent stochastic simulations started with initial conditions $n_{A}=n_{A}^{0}$ and $n_{i}=0$ ($i=1,\dots,S$) at $t=0$, each of which were stopped up to a time $t_{\text{end}}$ enough to reach stationarity. We maximized the log-likelihood function (75) using the L-BFGS-B optimization scheme (Zhu et al., 1997), which was used for numerical optimization throughout the contribution. Each realization of the data matrix $M(\{t_{j}\})$ leads to a line of maximum likelihood in the $(\alpha,\nu)$ space (see also Fig. 2, left panel), whose slope can vary over different replicas but can be determined precisely, providing to a consistent, unbiased estimation of the ratio $\widetilde{\nu}:=\nu/\alpha$ (Fig. 1, right panel). Using data sampled at stationarity, even if ML is based on a time-dependent log-likelihood, only parameter ratios are identifiable.
@@
From Eqs. (41)–(42), it is immediate to check that the dimensionless ratio $\widetilde{\nu}$ satisfies
$\widetilde{\nu}=\frac{\langle n_{A}\rangle}{\langle n_{AR}\rangle}\sum_{i=1}^{S}x_{R_{i}},$ (76)
@@
with $\langle n_{AR}\rangle:=\left\langle\sum_{i=1}^{S}n_{i}\right\rangle=\sum_{i=1}^{S}\langle n_{i}\rangle$ the average total number of handling consumers. In B we show that the likelihood function does exhibit a line of infinite critical points with slope given by (76), in which quantities $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are substituted by averages across samples, $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}$ and $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{S}n_{i,k}$, respectively. Therefore, our numerical experiment (Fig. 1 and Fig. (2), left panels) is consistent with a submanifold of degenerated log-likelihood optimum values in the parameter space.
@@
Figure 2: Heatmaps of negative log-likelihood. Representation of $-\log\mathcal{L}$, cf. Eq. (74), in the $(\alpha,\nu)$ space, obtained for two sets of $T=100$ model samples generated for $(\alpha,\nu)=(5,10)$: (i) taken at stationarity ($t_{\text{end}}\in[10\tau,20\tau]$, left panel), and (ii) taken out of equilibrium ($t_{\text{end}}\in[0,10\tau]$, right panel), $\tau$ being the characteristic time scale. Left: a line of infinite minima is obtained. Right: The arrow marks the true pair $(\alpha,\nu)=(5,10)$ used for data generation, and the white point marks the minimum of $-\log\mathcal{L}$. In both panels, remaining parameter values are the same as in Fig. 1.
@@
Figure 3: Distributions of parameter estimates. (a) The estimation procedure was repeated three times, for $n_{A}^{0}\in\{20,100,500\}$, using $T=20$ samples in simulated out-of-equilibrium data. Upper panels show the distributions of $\alpha^{\star}$ and $\nu^{\star}$, which tend to concentrate around true values (dashed lines) as $n_{A}^{0}$ increases. Parameter estimates show bias for low $n_{A}^{0}$ values, and the bias reduces as $n_{A}^{0}$ increase. This phenomenon is more clearly observed in the density plots represented in lower panels for $(\alpha^{\star},\nu^{\star})$, the lighter the color indicating the higher density of estimated pairs. (b) Same as (a) for $T=100$ temporal samples. Compared to panel (a), we observe that variability is decreased as $T$ increases. In both panels, remaining parameters are the same as in Fig. 1. $3000$ estimation experiments were conducted to obtain distributions.
@@
ML estimation techniques produce proper estimates for individual rates $\alpha$ and $\nu$ when data sample is taken out of equilibrium (and, importantly, at least two sampling times are different; see B). During the transient regime, samples were taken at times $\{t_{j}\}$ drawn from a uniform distribution over the interval $[0,10\tau]$, $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ being the characteristic dynamical time scale. Figure 2 compares the log-likelihood landscapes obtained using steady-state (left panel) and out-of-equilibrium model samples. A parameter screening in the $(\alpha,\nu)$ plane shows a line of infinite maxima $\nu^{\star}\propto\alpha^{\star}$ at equilibrium, confirming the results reported in Fig. 1. This evidences parameter redundancy when samples are taken at equilibrium. On the contrary, for out-of-equilibrium sampling, a single log-likelihood maximum value arises and parameters can be truly inferred from data. The accuracy of the estimates $\alpha^{\star}$ and $\nu^{\star}$ depends on data quality. For example, if the number of samples $T$ taken is large, then it is expected that estimates will be close to actual values. Similarly, if the overall number of individuals $n_{A}^{0}$ monitored in the experiment is large, then better estimates for model parameters are expected.
@@
Table 1: Confidence intervals. Here we present true values and $95\%$ confidence intervals for the in silico experiments shown in Fig. 3. The empirical distributions of parameter estimates, shown in that figure, are used to compute the corresponding percentiles that form the confidence interval. The sample of estimated parameters used to compute intervals has a size of $3000$ estimates for each $(T,n_{A}^{0})$ combination.
@@
Figure 4: Negative log-likelihood profiles for the same synthetic matrix $M(\{t_{j}\})$ used for the heatmap in Fig. 2 (right panel). Negative log-likelihood profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ (left and right panels, respectively) yield confidence intervals as the abscissae of the intercept of the profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ with the horizontal dotted lines whose ordinates are equal to $\ell_{1}(\alpha^{\star})+\chi^{2}_{1,0.95}/2$ and $\ell_{2}(\nu^{\star})+\chi^{2}_{1,0.95}/2$, respectively. $(\alpha^{\star},\nu^{\star})$ refers to the point (marked with vertical, dotted lines) at which the maximum of Eq. (74) is reached.
@@
$p_{A}(t)=\frac{\nu+\alpha\,e^{-\left(\nu+\alpha\sum_{k=1}^{S}x_{R_{k}}\right)t}\sum_{i=1}^{S}x_{R_{i}}}{\nu+\alpha\sum_{k=1}^{S}x_{R_{k}}},$ (77)
@@
Figure 5: Heatmaps of negative log-likelihood using only free predator abundances. Representation of $-\log\mathcal{L}$, cf. Eq. (78) (i.e., using only free predators abundance information), in the $(\alpha,\nu)$ space, obtained for the same two sets of model samples generated for Fig. 2: true parameter values are $(\alpha,\nu)=(5,10)$ (arrow in the right panel), and stationary (out-of-equilibrium) samples yield the log-likelihood represented in the left (right) panel. Observe that the heatmaps only differ from those in Fig. 2 in the colorbar values, because here we are using only partial information to evaluate $-\log\mathcal{L}$ for the same model observations as in that figure. Again, we find parameter redundancy for stationary sampling, whereas a single minimum is obtained for out-of-equilibrium sampling (the white point marks the minimum of $-\log\mathcal{L}$). Remaining parameter values were chosen as in Fig. 1.
@@
Figure 6: Distributions of parameter estimates using only free predator abundances. Same as Fig. 3 for a ML estimation procedure based on searching predator numbers. In this case, we have only used the number of free predators recorded at each sampling time, see Eq. (79). Distributions are quite similar to those obtained by using the vector of population handling predator abundances, $\bm{n}=(n_{1},\dots,n_{S})$, which are shown in Fig. 3. In both panels, parameters are the same as in Fig. 1.
@@
Table 2: Confidence intervals using only free predators abundances. The table lists actual values and $95\%$ confidence intervals for parameters of the in silico experiments shown in Fig. 6. Again, $3000$ samples of estimated parameters were used to compute intervals for each $(T,n_{A}^{0})$ combination.

### v3

@@ 3.1 Out-of-equilibrium probability distribution
This result agrees with those obtained at equilibrium, in the limit $t\to\infty$. As explained in A, we also found a multinomial distribution in the stationary regime. However, the out-of-equilibrium distribution for $\nu_{1}=\nu_{2}=\dots=\nu_{S}$ is not multinomial in general: note that the generating function for non-zero initial populations, Eq. (27), does not match the one one associated to a multinomial distribution, which is given by Eq. (28), unless there are no handling predators initially. $P_{\bm{n}}(t)$ follows a multinomial distribution only if $\bm{\hat{n}}=\bm{0}$. Irrespective of the set of initial conditions, however, the generating function (27) has the correct limit as $t\to\infty$. Indeed, if we take this limit on both sides of Eq. (27) we get
$\lim_{t\to\infty}G(\bm{z},t)=\left(1-\frac{m({\bm{z}})}{\nu+Q}\right)^{n_{A}^{0}}=\biggl{(}p_{A}+\sum_{j=1}^{S}p_{j}z_{j}\biggr{)}^{n_{A}^{0}},$
@@ 4 Effect of steady-state and out-of-equilibrium sampling
We replicate the experiment a number of times $T$, each replica simulating the stochastic dynamics (using Gillespie (1977) method) with $n_{A}^{0}$ free predators initially, and record abundances at time $t_{j}$ for each realization $j=1,\dots,T$. As shown in C, it is important that at least two sampling times are different (i.e., $t_{j}\neq t_{k}$ for some $j\neq k$). Otherwise the model is not identifiable. In practice, from an empirical point of view, it is very unlikely that independent system samples are obtained at precisely the same exact time.
@@
$M(\{t_{j}\})=\begin{pmatrix}n_{1,1}(t_{1})&n_{1,2}(t_{2})&\cdots&n_{1,T}(t_{T})\\ \vdots&\vdots&\ddots&\vdots\\ n_{S,1}(t_{1})&n_{S,2}(t_{2})&\cdots&n_{S,T}(t_{T})\end{pmatrix},$
where columns comprise vectors $\bm{n}_{j}=(n_{1,j},\dots,n_{S,j})$ reporting observed abundances of handling predators at sampling time $t_{j}$. Free predator abundance at time $t_{j}$ is computed with $n_{A}(t_{j})=n_{A}^{0}-\sum_{i=1}^{S}n_{i,j}(t_{j})$. If each column of $M(\{t_{j}\})$ is a sample of the Holling type II multi-resource dynamics, then the probability $P_{\bm{n}_{j}}(t_{j}|\bm{\theta})$, computed using Eqs. (11) and (29), is the probability of observing a population vector $(n_{A}(t_{j}),\bm{n}_{j})$ given a model parameter vector, $\bm{\theta}:=(\alpha,\nu)$ —notice that resource densities, $x_{R_{i}}=n_{R_{i}}/N$, the system volume, $N$, and the total number of consumers, $n_{A}^{0}$, are kept constant by assumption and therefore are not regarded as parameters in $\bm{\theta}$. Matrix $M(\{t_{j}\})$ gathers $T$ independent observations, so the function that assigns a likelihood to parameter vector $\bm{\theta}$ given the data matrix $M(\{t_{j}\})$ is multiplicative, $\mathcal{L}(\bm{\theta}|M(\{t_{j}\}))=\prod_{j=1}^{T}P_{\bm{n}_{j}}(t_{j}|\bm{\theta})$. The log-likelihood function,
@@
Figure 1: Estimation using equilibrium samples reveals parameter redundancy. We generated a single realization of the data matrix $M(\{t_{j}\})$, including $T=100$ independent model realizations starting from $n_{A}^{0}=100$ searching predators at $t=0$, using Gillespie’s stochastic simulation algorithm up to $T=100$ different ending times uniformly drawn in the interval $[50\tau,100\tau]$ to ensure stationarity, with $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ the characteristic time of the dynamics. True rates are $\alpha=5$, $\nu=10$, and individual consumers feed on $S=5$ resources with densities $\bm{x_{R}}=(0.9,0.7,0.5,0.3,0.1)$. Left: for that synthetic dataset, we run a L-BFGS-B optimization algorithm for the log-likelihood function (30) starting from different initial conditions, and we recorded the values of $(\alpha^{\star},\nu^{\star})$ at which the maximum was reached. Optimal parameter values follow a perfect straight line. The distributions of $\alpha^{\star}$ and $\nu^{\star}$ are compatible with uniform distributions, and the optimal values of $\log\mathcal{L}$ keep constant up to round-off errors. Overall, this indicates a line of infinite optima. Right: we replicated the previous experiment 200 times and, for each replica, we estimated the ratio $\nu/\alpha$ by fitting the slope of the straight line. We obtained the distribution in the right panel, which is compatible with a Gaussian with mean 2 and variance extracted from data. The vertical, dashed line indicates the true ratio.
@@
Now, consider the effect of sampling times and assume that matrix $M(\{t_{j}\})$ is obtained when $t_{j}$ are “large enough” compared to the characteristic time of the dynamics, which is given by $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ —see Eq. (29). In Fig. 1, in silico data (i.e., abundances of each behavioral type) were generated by stochastic simulation of a system with $S$ resources (with fixed densities $x_{R_{i}}$) consumed by $n_{A}^{0}$ predator individuals. A number $T$ of independent stochastic simulations started with initial conditions $n_{A}=n_{A}^{0}$ and $n_{i}=0$ ($i=1,\dots,S$) at $t=0$, each of which were stopped up to a time $t_{j}$ drawn uniformly in $[50\tau,100\tau]$, which is enough to reach stationarity. We maximized the log-likelihood function (30) using the L-BFGS-B optimization scheme (Zhu et al., 1997), which was used for numerical optimization throughout the contribution. Each realization of the data matrix $M(\{t_{j}\})$ leads to a line of maximum likelihood in the $(\alpha,\nu)$ space (see also Fig. 2, left panel), whose slope can vary over different replicas but can be determined precisely, providing to a consistent, unbiased estimation of the ratio $\nu/\alpha$ (Fig. 1, right panel). Using data sampled at stationarity, even if ML is based on a time-dependent log-likelihood, only parameter ratios are identifiable.
@@
From Eqs. (46)–(47) in A, it is immediate to check that the dimensionless ratio $\nu/\alpha$ satisfies
$\frac{\nu}{\alpha}=\frac{\langle n_{A}\rangle}{\langle n_{AR}\rangle}\sum_{i=1}^{S}x_{R_{i}},$ (31)
@@
with $\langle n_{AR}\rangle:=\left\langle\sum_{i=1}^{S}n_{i}\right\rangle=\sum_{i=1}^{S}\langle n_{i}\rangle$ the average total number of handling consumers. In C we show that, when sampling times coincide ($t_{1}=t_{2}=\dots=t_{T}=t$) and $t\to\infty$, the likelihood function does exhibit a line of infinite critical points with slope given by (31), in which quantities $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are substituted by averages across samples, $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}$ and $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{S}n_{i,k}$, respectively. Therefore, our numerical experiment (Fig. 1 and Fig. (2), left panels) is consistent with a submanifold of degenerated log-likelihood optimum values in the parameter space.
@@
Figure 2: Heatmaps of negative log-likelihood for stationary and out-of-equilibrium samples unveils redundancy (for steady-state sampling). Representation of $-\log\mathcal{L}$, cf. Eq. (30), in the $(\alpha,\nu)$ space, obtained for two sets of $T=100$ model samples generated for $(\alpha,\nu)=(5,10)$: (i) taken at stationarity (sampling times $\{t_{j}\}$ drawn uniformly in $[10\tau,20\tau]$, left panel), and (ii) taken out of equilibrium (sampling times $\{t_{j}\}$ drawn uniformly in $[0,10\tau]$, right panel), $\tau$ being the characteristic time scale. Left: a line of infinite minima is obtained. Right: The arrow marks the true pair $(\alpha,\nu)=(5,10)$ used for data generation, and the white point marks the minimum of $-\log\mathcal{L}$. In both panels, remaining parameter values are the same as in Fig. 1.
@@
Figure 3: Distributions of ML parameter estimates are unbiased for different number of initial consumers and increased number of temporal samples. (a) The estimation procedure was repeated three times, for $n_{A}^{0}\in\{20,100,500\}$, using $T=20$ samples in simulated out-of-equilibrium data. Upper panels show the distributions of $\alpha^{\star}$ and $\nu^{\star}$, which tend to concentrate around true values (dashed lines) as $n_{A}^{0}$ increases. Parameter estimates show bias for low $n_{A}^{0}$ values, and the bias reduces as $n_{A}^{0}$ increase. This phenomenon is more clearly observed in the density plots represented in lower panels for $(\alpha^{\star},\nu^{\star})$, the lighter the color indicating the higher density of estimated pairs. (b) Same as (a) for $T=100$ temporal samples. Compared to panel (a), we observe that variability is decreased as $T$ increases. In both panels, remaining parameters are the same as in Fig. 1. $3000$ estimation experiments were conducted to obtain distributions.
@@
ML estimation techniques produce proper estimates for individual rates $\alpha$ and $\nu$ when data sample is taken out of equilibrium (and, importantly, at least two sampling times are different; see C). During the transient regime, samples were taken at times $\{t_{j}\}$ drawn from a uniform distribution over the interval $[0,10\tau]$, $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ being the characteristic dynamical time scale. Figure 2 compares the log-likelihood landscapes obtained using steady-state (left panel) and out-of-equilibrium model samples (taken at times uniformly drawn over the interval $[10\tau,20\tau]$). A parameter screening in the $(\alpha,\nu)$ plane shows a line of infinite maxima $\nu^{\star}\propto\alpha^{\star}$ at equilibrium, confirming the results reported in Fig. 1. This evidences parameter redundancy when out-of-equilibrium samples are missing. On the contrary, a single log-likelihood maximum value arises and parameters can be truly inferred from the data when out-of-equilibrium samples are included (we have checked that, if samples are taken at times within the interval $[0,20\tau]$, we also find a single maximum in log-likelihood, which implies that redundancy is avoided even when the data contain steady-state and out-of-equilibrium observations together). The accuracy of the estimates $\alpha^{\star}$ and $\nu^{\star}$ depends on data quality. For example, if the number of samples $T$ taken is large, then it is expected that estimates will be close to actual values. Similarly, if the overall number of individuals $n_{A}^{0}$ monitored in the experiment is large, then better estimates for model parameters are expected.
@@
Table 1: Confidence intervals calculated using parameter distributions. Here we present true values and $95\%$ confidence intervals for the in silico experiments shown in Fig. 3. The empirical distributions of parameter estimates, shown in that figure, are used to compute the corresponding percentiles that form the confidence interval. The sample of estimated parameters used to compute intervals has a size of $3000$ estimates for each $(T,n_{A}^{0})$ combination.
@@
Figure 4: Confidence interval estimation using negative log-likelihood profiles for the same synthetic matrix $M(\{t_{j}\})$ used for the heatmap in Fig. 2 (right panel). Negative log-likelihood profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ (left and right panels, respectively) yield confidence intervals as the abscissae of the intercept of the profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ with the horizontal dotted lines whose ordinates are equal to $\ell_{1}(\alpha^{\star})+\chi^{2}_{1,0.95}/2$ and $\ell_{2}(\nu^{\star})+\chi^{2}_{1,0.95}/2$, respectively. $(\alpha^{\star},\nu^{\star})$ refers to the point (marked with vertical, dotted lines) at which the maximum of negative log-likelihood is reached.
@@
$p_{A}(t)=\frac{\nu+\alpha\,e^{-\left(\nu+\alpha\sum_{k=1}^{S}x_{R_{k}}\right)t}\sum_{i=1}^{S}x_{R_{i}}}{\nu+\alpha\sum_{k=1}^{S}x_{R_{k}}},$
@@
Figure 5: Heatmaps of negative log-likelihood using only free predator abundances for stationary and out-of-equilibrium samples unveils redundancy (for steady-state sampling). Representation of $-\log\mathcal{L}$, cf. Eq. (32) (i.e., using only free predators abundance information), in the $(\alpha,\nu)$ space, obtained for the same two sets of model samples generated for Fig. 2: true parameter values are $(\alpha,\nu)=(5,10)$ (arrow in the right panel), and stationary (out-of-equilibrium) samples yield the log-likelihood represented in the left (right) panel. Observe that the heatmaps only differ from those in Fig. 2 in the colorbar values, because here we are using only partial information to evaluate $-\log\mathcal{L}$ for the same model observations as in that figure. Again, we find parameter redundancy for stationary sampling, whereas a single minimum is obtained for out-of-equilibrium sampling (the white point marks the minimum of $-\log\mathcal{L}$). Remaining parameter values were chosen as in Fig. 1.
@@
Figure 6: Distributions of ML parameter estimates using only free predator abundances are unbiased for increasing number of initial consumers and number of temporal samples. Same as Fig. 3 for a ML estimation procedure based on searching predator numbers. In this case, we have only used the number of free predators recorded at each sampling time, see Eq. (4). Distributions are quite similar to those obtained by using the vector of population handling predator abundances, $\bm{n}=(n_{1},\dots,n_{S})$, which are shown in Fig. 3. In both panels, parameters are the same as in Fig. 1.
@@
Table 2: Confidence intervals using only free predators abundances calculated using parameter distributions (Fig. 6). The table lists actual values and $95\%$ confidence intervals for parameters of the in silico experiments shown in Fig. 6. Again, $3000$ samples of estimated parameters were used to compute intervals for each $(T,n_{A}^{0})$ combination.